# Exponentially Weighted Moving Models

### v1

@@ 1.1 Previous and related work
The idea of the exponentially weighted moving average (EWMA) is well known, and has origins going as far back as the recursive exponential window functions used by Poisson in the 19th century. It was introduced to statistics in 1956 by Brown as a method for forecasting demand [Bro56]. In the context of signal processing, the EWMA is an application of a window function, used as a low pass filter to remove high frequency noise from a signal [OSB99]. Moving averages are related to the concept of rolling window estimation, wherein models are repeatedly fit to a fixed size window of past data. RWMs are widely used in time series analysis and forecasting in economics, finance, and engineering [Box+15, Tsa02].
Exponentially weighted sums also appear in finance, albeit applied to future rather than past terms. Cash flows are discounted in the calculation of present value; here the discount factor is $\frac{1}{1+r}$, where $r$ is the interest rate. Exponential weighting of future terms also appears in Markov decision processes (MDPs) [Ber22], where the value function represents the discounted sum of future rewards. Exponentially weighted moving averages appear in the context of model free reinforcement learning, where an exponentially decayed expectation of rewards is incrementally estimated via a recursive update [KWW22, §17.1-3].
@@
In online quantile estimation, the goal is to estimate the $\eta$-quantile of a time series, for some fixed $\eta\in(0,1)$. This can be computed exactly at any time by storing all past data and sorting it, but this need not be practical for large data sets. Several methods exist for online quantile estimation, such as the $P^{2}$ algorithm for online quantile estimation without storing all points [JC85]. This method stores only 5 points chosen on the empirical CDF. Another well known work by Greenwald and Khanna [GK01] provides a space efficient method based on defining a notion of approximate quantiles, where the approximation error is allowed to grow with the number of data points.
@@
The EWMM is a generalization of the exponentially weighted moving regression model, which has been studied in the context of time series forecasting. The original work on this topic is due to [Chr71], who extended the method described by Brown [Bro56] to the case of linear regression, which allowed for the use of features in the forecasting model. Another related area of work is locally weighted regression, which also fits a regression model with a weighted sum of error functions [Cle79]. The weights are typically chosen to be a function of the distance between the point at which a prediction is being made and the data points being used to fit the model. In our setting, the notion of distance is temporal rather than spatial, and the weight decreases exponentially with distance, i.e., lapsed time. Hastie and Tibshirani [TH87] provide a treatment of local likelihood estimation, which generalizes moving window linear regression to likelihood based regression models, where maximum likelihood estimates are computed on a window of data points near the point at which a prediction is being made. In the case of moving window linear regression, a classic algebraic trick that allows for reduced computational complexity of the recursive update of estimates has been known since Gauss and Legendre [Sor70, GS95].
@@
In computer science, online learning concerns the problem of learning from a stream of data, where the goal is to make predictions about the next data point, and update the model based on the observed data. This is in contrast to batch learning, where the model is trained on a fixed set of data. See [McM17] for a survey of online learning algorithms. Hazan provides a comprehensive overview of online convex optimization in [Haz16]. The exponentially weighted moving model can be considered a form of online learning. However, much of the work in online learning focuses on the regret, the difference between the performance of the online learning algorithm and the best fixed model in hindsight. The motivation for EWMM differs in that our goal is to estimate a time-varying parameter well at each time period. But like EWMM, online learning solves a problem, typically convex and of fixed size, each time period to update its estimates of the parameters.
@@
The idea of approximating part of an objective function as a convex quadratic is a basic one in several optimization methods, most famously Newton’s method [BV04, §9.5]. A sophisticated extension is sequential quadratic programming (SQP) methods [BT95]. In the context of control, quadratic approximations of the value function lead to approximate dynamic programming (ADP methods) [Pow11, KB14]. A convex quadratic terminal cost or value function in often used in model predictive control (MPC) [WB09, WOB15].
@@ 2.3.1 Exponentially weighted moving data models
Instead of quadratic loss we can use a robust loss function such as the Huber loss, which would give the exponentially weighted moving robust estimate of the mean [Hub92].
@@
where $\eta\in[0,1]$ is the quantile level, we obtain the exponentially weighted moving estimate of the $\eta$-quantile [KBJ78].
@@ 2.3.2 Exponentially weighted moving regression models
We use loss $\ell(x,\theta)=L(y_{t}-\hat{y}_{t})$, where $L$ is a convex loss function. With $L(u)=\|u\|_{2}^{2}$, we get the exponentially weighted ordinary least squares regression model. We can use other losses such as pinball or Huber. We can add any convex regularization. With regularizer $r(\theta)=\lambda\|\theta\|_{2}^{2}$, where $\lambda>0$ is a hyper-parameter, we obtain exponentially weighted ridge regression [GVL89, page 564] With $r(\theta)=\lambda\|\theta\|_{1}$, we obtain the exponentially weighted LASSO regression model [Tib96]. With $r(\theta)=0$ for $\theta\geq 0$ (elementwise) and $r(\theta)=\infty$ otherwise, we obtain the exponentially weighted nonnegative least squares regression model.
@@
we obtain exponentially weighted logistic regression [HTF01].
@@ 3.2 Examples
with $\lambda>0$ [FHT07]. We recursively compute the EWMA empirical estimate
@@
to obtain an exponentially weighted PMF estimate that is smooth with respect to the graph [TB21].
@@
where $e_{j}$ is the standard $j$th unit vector in ${\mbox{\bf R}}^{m}$, i.e., $(u_{j})_{k}=1$ if $k=j$ and $(u_{j})_{k}=0$ if $k\neq j$. The first term on the righthand side is linear in $\theta$, and therefore also quadratic, do we take that as our loss. We take the second and third terms on the righthand side as the regularizer. The vector in parentheses in the first term on the righthand side is the EWMA estimate of the past frequencies of occurrence, which of course can be computed recursively.
@@
The fact that the EWMM for exponential families can be computed with a finite size problem is connected to a well known result in statistics, the Pitman-Koopman-Darmois theorem [Pit36, Koo36, Dar35]. The theorem says that, under some minor technical conditions, the exponential family of distributions is the only family where there can be a sufficient statistic whose dimension does not grow with the sample size.
@@ 4.3 Tail fitting
There are many principled methods to generate the points $u_{i}$. If the dimension of $\theta$ is small enough, we can choose a uniform grid of points around $\theta_{t-1}$. One could also use a low discrepancy sequence [Sob67] as a more sophisticated way to cover the space. One can also model the variance of $\theta$ across previous estimates and generate points using so-called sigma points [VDMW04]. Even simpler is to sample from a normal distribution centered at $\theta_{t-1}$. See [KW19, Chap. 13] for a detailed study of methods for generating evaluation points for approximation problems. Any method of generating the evaluation points should exclude any points not in $\Theta$.
@@ 5 Numerical examples
In this section we give some examples of evaluating the EWMM either exactly (our first example) or approximately (for the others). All examples can be reproduced using publicly available code at https://github.com/cvxgrp/ewmm_code. We use CVXPY, a Python-embedded modeling language for convex optimization to specify and compute the EWMM [DB16].
@@ 5.1 Sparse inverse covariance estimation
We use the 10 Industry Portfolio dataset from Kenneth French’s data library [Fre24]. The dataset contains daily returns for 10 value-weighted industry portfolios. We examine returns from the last 4 years, from 2020-01-02 to 2024-01-31 giving us data $x_{1},\ldots,x_{T}\in{\mbox{\bf R}}^{10}$, where $T=1027$.

### v2

@@ 1.1 Previous and related work
The idea of the exponentially weighted moving average (EWMA) is well known, and has origins going as far back as the recursive exponential window functions used by Poisson in the 19th century. It was introduced to statistics in 1956 by Brown as a method for forecasting demand [adl1956exponential]. In the context of signal processing, the EWMA is an application of a window function, used as a low pass filter to remove high frequency noise from a signal [oppenheim1999discrete]. Moving averages are related to the concept of rolling window estimation, wherein models are repeatedly fit to a fixed size window of past data. RWMs are widely used in time series analysis and forecasting in economics, finance, and engineering [box2015time, tsay2005analysis].
Exponentially weighted sums also appear in finance, albeit applied to future rather than past terms. Cash flows are discounted in the calculation of present value; here the discount factor is $\frac{1}{1+r}$, where $r$ is the interest rate. Exponential weighting of future terms also appears in Markov decision processes (MDPs) [bertsekas2022abstract], where the value function represents the discounted sum of future rewards. Exponentially weighted moving averages appear in the context of model free reinforcement learning, where an exponentially decayed expectation of rewards is incrementally estimated via a recursive update [kochenderfer2022algorithms, §17.1-3].
@@
In online quantile estimation, the goal is to estimate the $\eta$-quantile of a time series, for some fixed $\eta\in(0,1)$. This can be computed exactly at any time by storing all past data and sorting it, but this need not be practical for large data sets. Several methods exist for online quantile estimation, such as the $P^{2}$ algorithm for online quantile estimation without storing all points [jain1985p2]. This method stores only 5 points chosen on the empirical CDF. Another well known work by Greenwald and Khanna [greenwald2001space] provides a space efficient method based on defining a notion of approximate quantiles, where the approximation error is allowed to grow with the number of data points.
@@
The EWMM is a generalization of the exponentially weighted moving regression model, which has been studied in the context of time series forecasting. The original work on this topic is due to [christiaanse1971short], who extended the method described by Brown [adl1956exponential] to the case of linear regression, which allowed for the use of features in the forecasting model. Another related area of work is locally weighted regression, which also fits a regression model with a weighted sum of error functions [cleveland1979robust]. The weights are typically chosen to be a function of the distance between the point at which a prediction is being made and the data points being used to fit the model. In our setting, the notion of distance is temporal rather than spatial, and the weight decreases exponentially with distance, i.e., lapsed time. Hastie and Tibshirani [tibshirani1987local] provide a treatment of local likelihood estimation, which generalizes moving window linear regression to likelihood based regression models, where maximum likelihood estimates are computed on a window of data points near the point at which a prediction is being made. In the case of moving window linear regression, a classic algebraic trick that allows for reduced computational complexity of the recursive update of estimates has been known since Gauss and Legendre [sorenson1970least, gauss1995theory].
@@
In computer science, online learning concerns the problem of learning from a stream of data, where the goal is to make predictions about the next data point, and update the model based on the observed data. This is in contrast to batch learning, where the model is trained on a fixed set of data. See [mcmahan2017survey] for a survey of online learning algorithms. Hazan provides a comprehensive overview of online convex optimization in [hazan2016introduction]. The exponentially weighted moving model can be considered a form of online learning. However, much of the work in online learning focuses on the regret, the difference between the performance of the online learning algorithm and the best fixed model in hindsight. The motivation for EWMM differs in that our goal is to estimate a time-varying parameter well at each time period. But like EWMM, online learning solves a problem, typically convex and of fixed size, each time period to update its estimates of the parameters.
@@
The idea of approximating part of an objective function as a convex quadratic is a basic one in several optimization methods, most famously Newton’s method [cvxbook, §9.5]. A sophisticated extension is sequential quadratic programming (SQP) methods [boggs1995sequential]. In the context of control, quadratic approximations of the value function lead to approximate dynamic programming (ADP methods) [powell2007approximate, keshavarz2014quadratic]. A convex quadratic terminal cost or value function in often used in model predictive control (MPC) [wang2009performance, wang2015approximate].
@@ 2.3.1 Exponentially weighted moving data models
Instead of quadratic loss we can use a robust loss function such as the Huber loss, which would give the exponentially weighted moving robust estimate of the mean [huber1992robust].
@@
where $\eta\in[0,1]$ is the quantile level, we obtain the exponentially weighted moving estimate of the $\eta$-quantile [koenker1978regression].
@@ 2.3.2 Exponentially weighted moving regression models
We use loss $\ell(x,\theta)=L(y_{t}-\hat{y}_{t})$, where $L$ is a convex loss function. With $L(u)=\|u\|_{2}^{2}$, we get the exponentially weighted ordinary least squares regression model. We can use other losses such as pinball or Huber. We can add any convex regularization. With regularizer $r(\theta)=\lambda\|\theta\|_{2}^{2}$, where $\lambda>0$ is a hyper-parameter, we obtain exponentially weighted ridge regression [golub2013matrix, page 564] With $r(\theta)=\lambda\|\theta\|_{1}$, we obtain the exponentially weighted LASSO regression model [tibshirani1996regression]. With $r(\theta)=0$ for $\theta\geq 0$ (elementwise) and $r(\theta)=\infty$ otherwise, we obtain the exponentially weighted nonnegative least squares regression model.
@@
we obtain exponentially weighted logistic regression [hastie2009elements].
@@ 3.2 Examples
with $\lambda>0$ [friedman2007sparse]. We recursively compute the EWMA empirical estimate
@@
to obtain an exponentially weighted PMF estimate that is smooth with respect to the graph [tuck2021fitting].
@@
where $e_{j}$ is the standard $j$th unit vector in ${\mbox{\bf R}}^{m}$, i.e., $(e_{j})_{k}=1$ if $k=j$ and $(e_{j})_{k}=0$ if $k\neq j$. The first term on the righthand side is linear in $\theta$, and therefore also quadratic, do we take that as our loss. We take the second and third terms on the righthand side as the regularizer. The vector in parentheses in the first term on the righthand side is the EWMA estimate of the past frequencies of occurrence, which of course can be computed recursively.
@@
The fact that the EWMM for exponential families can be computed with a finite size problem is connected to a well known result in statistics, the Pitman-Koopman-Darmois theorem [pitman1936sufficient, koopman1936distributions, darmois1935lois]. The theorem says that, under some minor technical conditions, the exponential family of distributions is the only family where there can be a sufficient statistic whose dimension does not grow with the sample size.
@@ 4.3 Tail fitting
There are many principled methods to generate the points $u_{i}$. If the dimension of $\theta$ is small enough, we can choose a uniform grid of points around $\theta_{t-1}$. One could also use a low discrepancy sequence [sobol1967distribution] as a more sophisticated way to cover the space. One can also model the variance of $\theta$ across previous estimates and generate points using so-called sigma points [van2004sigma]. Even simpler is to sample from a normal distribution centered at $\theta_{t-1}$. See [kochenderfer2019algorithms, Chap. 13] for a detailed study of methods for generating evaluation points for approximation problems. Any method of generating the evaluation points should exclude any points not in $\Theta$.
@@ 5 Numerical examples
In this section we give some examples of evaluating the EWMM either exactly (our first example) or approximately (for the others). All examples can be reproduced using publicly available code at https://github.com/cvxgrp/ewmm_code. We use CVXPY, a Python-embedded modeling language for convex optimization to specify and compute the EWMM [diamond2016cvxpy].
@@ 5.1 Sparse inverse covariance estimation
We use the 10 Industry Portfolio dataset from Kenneth French’s data library [frenchdata]. The dataset contains daily returns for 10 value-weighted industry portfolios. We examine returns from the last 4 years, from 2020-01-02 to 2024-01-31 giving us data $x_{1},\ldots,x_{T}\in{\mbox{\bf R}}^{10}$, where $T=1027$.